The following is a list of accounts of the Dienstag Company for the current year end (December 31, 2020. All of the accountsCreate a statement of Retained earnings
Answer #1

DIENSTAG COMPANY

Statement of Retained earnings

Fot the year ending december31,2020

Retained earning on january1,2020 568000
add: net income earned in 2020 28800
total 596800
less: dividend declared (120000)
reatined earning 476800

* net income = (revenue - expenses) - income tax

revenue : sales - cost of goods sold , 16000000- (560000+440000-32000)=1072000

total expenses = 1024000

income before tax = 48000

net income = 48000-40% = 28800